If (an) is convergent then "\sum_{i=1}^{\infty} a_n" is also convergent. True or false with the full explanation.


1
Expert's answer
2022-06-10T09:21:20-0400

The sequence "a_n=\dfrac{1}{n}" converges to "0."

But the harmonic series


"\displaystyle\sum_{i=1}^{\infin}\dfrac{1}{n}"

is divergent.

This is the counterexample of the statement.

Therefore the statement is False.